Solve the quadratic equation by factoring and applying the property "ab=0 if and only if a=0 or b=0."
(x^2)+4kx=0.

Answer by stanbon(75887)

(x^2)+4kx=0.
Factor to get:
x(x+4k)=0
x=0 or x+4k=0
x=0 or x=-4k
Cheers,
Stan H.